OpenStudy (anonymous):

The surface area, S, of a cylinder is given by the formula s=2PIrh+2PIr^2. Which equation is equivalent to s=2PIrh+2PIr^2? a) h=s-2PIr-2PIr^2 b) h=s+2PIr+2PIr^2 c) h=s/2PIr-2PIr^2 d) h=s/2PIr-r

OpenStudy (jdoe0001):

let us say... you have a = b + c what would be your first step to solve for "b"?

OpenStudy (anonymous):

S=2*pi*r*h+2*pi*r^2 subtract 2*pi*r^2 S-2*pi*r^2=2*pi*r*h divide by 2*pi*r to leave h (S-2*pi*r^2)/2*pi*r so there is a mistake in your question

OpenStudy (anonymous):

Thank you @jdoe0001 and @jamaica25